# Generalizing Graph Transformers Across Diverse Graphs and Tasks via Pre-Training on Industrial-Scale Data

### v3

@@ 4.2. Masked Encoding
Differing from most existing masked autoencoder-based graph self-supervised learning frameworks that employ GNNs as encoders (Hou et al., 2022; Li et al., 2023), in this work, we use the Transformer model as the encoder, which introduces differences in the pipeline.
@@
For a seed node $v_{s}$ in the graph, we sample a sequence of contextual nodes $S_{s}=(v_{s},v_{s,1},v_{s,2},\ldots,v_{s,k})$ of length $k+1$ including seed node $v_{s}$ itself. A subset of this contextual sequence is randomly masked, resulting in a masked set $S^{m}_{s}=\{v_{s,m_{1}},v_{s,m_{2}},\ldots,v_{s,m_{l}}\}$ of length $l$, where $l<k+1$. The remaining unmasked nodes are represented as $S^{u}_{s}=\{v_{s,u_{1}},v_{s,u_{2}},\ldots,v_{s,u_{k-l+1}}\}$. The node feature $\mathbf{x}_{s,u_{j}}\in\mathbb{R}^{1\times d}$ of each unmasked node $v_{s,u_{j}}$ is first mapped into an embedding space using a linear projection from $d$ dimension to $d^{\prime}$ dimension, where $10<j\leq k-l+$ 1. In addition, we add pre-computed positional encodings to the node embedding:
(1) $\mathbf{h}_{s,u_{j}}=\mathrm{Linear}(\mathbf{x}_{s,u_{j}})+\mathbf{p}_{s,u_{j}}$
These embeddings serve as the input to the transformer. The embedding matrix of unmasked node sequence $\mathbf{H}^{u}_{s}\in\mathbb{R}^{(k-l+1)\times d^{\prime}}$ are passed through a $L$-layer transformer encoder. In each layer, the transformer computes self-attention weights and aggregates information from all nodes in the sequence. The output after the final encoder layer is $\mathbf{H}^{u}_{s}$, which represents the hidden representation of unmasked node sequence $S^{u}_{s}$ within the transformer. The specific operations in a transformer layer can be described as follows:$\mathbf{H}^{u}_{s}=f_{E}(\mathbf{H}^{u}_{s})$, where $f_{E}(\cdot)$ represents the operations of the encoder of PGT, which is a $L$-layers transformer encoder.
4.3. Pre-training Tasks
@@
Feature reconstruction. Given a node sequence $S_{s}$ along with its corresponding node features $\mathbf{X}_{s}\in\mathbb{R}^{(k+1)\times d}$, after the masked encoding phase, we obtain representations $\mathbf{H}^{u}_{s}\in\mathbb{R}^{(k-l+1)\times d^{\prime}}$ for the unmasked node sequence. Similar to MIM (He et al., 2022), the input to the decoder is the full set of tokens consisting of encoded unmasked nodes and mask nodes. We employ a shared learnable token $\mathbf{h}_{[M]}\in\mathbb{R}^{1\times d^{\prime}}$ to initialize the representations of these masked nodes, and this learnable token is continually updated during training. At the same time, we once again add pre-computed positional encodings into the corresponding node embeddings to introduce graph structural information to the decoding phase. The input representation matrix $\mathbf{H}_{s}\in\mathbb{R}^{(k+1)\times d^{\prime}}$ of the decoder can be denoted as:
@@
Then the decoder would reconstruct the input features $\mathbf{X}_{s}$ from the latent embeddings $\mathbf{H}_{s}$. In this task, we use another relatively shallow transformer as a decoder, denoted as $f_{D_{1}}(\cdot)$. The reconstructed features are $\mathbf{Z}_{s}=f_{D_{1}}(\mathbf{H}_{s})$. We adopt scaled cosine error (Hou et al., 2022) as the objective function for feature reconstruction:
(2) $\mathcal{L}_{1}=\frac{1}{|S^{m}_{s}|}\sum_{v_{s,i}\in S^{m}_{s}}(1-\frac{\mathbf{x}_{i}^{\top}\mathbf{z}_{i}}{\lVert\mathbf{x}_{i}\lVert\cdot\lVert\mathbf{z}_{i}\lVert})^{\gamma}$
@@
Local graph structure reconstruction. Our intuition behind using PPR to sample contextual node sequences is that selecting the most relevant neighboring nodes can create small and densely connected clusters. PPR can be interpreted within the spectral framework as a localized spectral method (Green et al., 2021). Spectral clustering essentially operates in the eigenspace of the graph Laplacian matrix, capturing the low-frequency components that represent tightly-knit communities. (Tremblay and Loukas, 2020). Similarly, the embeddings learned by the proposed PGT should ideally reside in a subspace where the local structural properties are maintained. This task aims to capture local graph structure patterns by leveraging the likelihood that nodes within the same sequence exhibit higher similarity, in contrast to nodes from different sequences, which are generally less similar.
@@
$\mathcal{P}$ =$=\{(\mathbf{h}_{i,j},\mathbf{h}_{i,k}):i\in\{1,...N\},j,k\in\{1,...M_{i}\},j\neq k\},$
$\mathcal{N}$ =$=\{(\mathbf{h}_{i,j},\mathbf{h}_{l,m}):i,l\in\{1,...N\},i\neq l,$
$j\in\{1,...M_{i}\},m\in\{1,...M_{l}\}\}.$
@@
(3) $s_{a,b}=\text{sim}\left(f_{D_{2}}(\mathbf{h}_{a}),f_{D_{2}}(\mathbf{h}_{b})\right)$
@@
(4) $\mathcal{L}_{2}=-\frac{1}{T}\sum_{(\mathbf{h}_{i,j},\mathbf{h}_{i,k})\in\tilde{\mathcal{P}}}\log\left(\frac{\exp(s_{ij,ik})}{\sum_{(\mathbf{h}_{l,m},\mathbf{h}_{n,o})\in\tilde{\mathcal{P}}\cup\tilde{\mathcal{N}}}\exp(s_{lm,no})}\right).$
@@
(5) $\mathcal{L}=\mathcal{L}_{1}+\lambda\cdot\mathcal{L}_{2}$
4.4. Decoder Reuse for Feature Augmentation
Under the classic auto-encoder paradigm, only the pre-trained encoder is used in downstream tasks, and the decoder is discarded (He et al., 2022; Hou et al., 2022, 2023; Li et al., 2023). In PGT, we propose a novel approach to utilize the pre-trained decoder for feature augmentation when the pre-trained encoder is used for inference or finetuned for downstream tasks. Given a contextual node sequence $S_{q}$ of a query node $v_{q}$, we generate reconstructed node features by passing the node sequences and their features $\mathbf{X^{\prime}_{q}}\in\mathbb{R}^{(k+1)\times d}$ into the pre-trained encoder and decoder for features reconstruction without any mask:
(6) $\hat{\mathbf{H^{\prime}_{q}}}=f_{E}(\mathrm{Linear}(\mathbf{X^{\prime}_{q}})+\mathbf{P^{\prime}_{q}})$
(7) $\overline{\mathbf{X^{\prime}_{q}}}=f_{D_{1}}(\hat{\mathbf{H^{\prime}_{q}}}+\mathbf{P^{\prime}_{q}})$
@@ 4.4. Decoder Reuse for Feature Augmentation
Inference. The process of utilizing pre-trained decoder for feature augmentation is decoupled from the subsequent fine-tuning of encoder for downstream tasks. After obtaining the newly augmented features, decoders are discarded, and only the encoder is employed for generating node representations or subject to fine-tuning. We take the output of the last layer of encoder as node embeddings:
(8) $\mathbf{H^{\prime}_{q}}=f_{E}(\mathrm{Linear}(\frac{1}{2}(\mathbf{X^{\prime}_{q}}+\overline{\mathbf{X^{\prime}_{q}}}))+\mathbf{P^{\prime}_{q}})$
@@
5. Deployment in Tencent
5.1. Pre-training
Table 1. Statistics of datasets for pre-training and Minor Detection Task.
@@ 5.1. Pre-training
Table 2. Linear probing results on Minors Detection Task. We report ROC-AUC for all datasets. No-pretrain represents a random-initialized model without any self-supervised pretraining.
@@
Table 3. Results of fine-tuning the pre-trained model with 1%, 5% and 100% labeled training data on Minors Detection Task. We report ROC-AUC for all datasets. No-pretrain represents a random-initialized model without any self-supervised pre-training.
@@
Dataset. We clarify some fundamental concepts in the online gaming business. Tencent’s online games and social media platforms are closely interconnected. Players can log into various games using their social platform accounts, such as WeChat, and can add new game friends within the games. For our pre-training dataset, we select game players on the WeChat platform as nodes and the union of each player’s friend relationships in a few games as edges. For privacy reasons, we extract a relatively large subgraph from the raw data. All subsequently mentioned datasets are constructed in this manner. We construct a 146-dimensional numerical feature for each node. These features include information about the account’s login history, activity duration, and in-game item purchase records over a certain period. Detailed information is shown in Table 1.
Baselines. We compare our PGT framework with state-of-the-art graph pre-training and self-supervised learning algorithms, DGI (Veličković et al., 2018), GRACE (Zhu et al., 2020), BGRL (Thakoor et al., 2021), GraphMAE (Hou et al., 2022) and GraphMAE2 (Hou et al., 2023). Other methods are not included in the comparison because they are not scalable to large graphs, e.g., GMAE (Zhang et al., 2022a). Given the enormous scale of the pre-training graph, sampling algorithms like GraphSAINT (Zeng et al., 2019) or Neighbor Sampling (Hamilton et al., 2017) require more than 72 hours for pre-training. We adopt ClusterGCN (Chiang et al., 2019) as the subgraph sampling method in our baselines due to its superior training efficiency. For all baseline methods, we employ GAT (Veličković et al., 2017) as the backbone network.
5.2. Minor Detection Task
Task description. By relevant national regulations, Tencent’s online games have restrictions on the playing time of minors. The objective of the minors detection task is to identify users in a specific online game who appear to be minors, to take further measures. In this paper, we model it as a binary classification task per node on the graph. For each node to be predicted, we classify it as either an adult or a minor.
Setup. We built our datasets from the accounts of Game A and B, as illustrated in Table 1. For both datasets, the training, validation, and test sets consist of 100,000, 100,000, and 800,000 nodes, respectively.
We select all the self-supervised pre-trained models mentioned earlier as baseline methods. In alignment with pre-training, we employ ClusterGCN as the subgraph sampling algorithm to scale baseline methods to large graphs. Following most of the graph self-supervised learning algorithms (Hou et al., 2023), we evaluate our method under two settings: linear probing and fine-tuning. For the linear probing setting, we freeze the pre-trained encoder and infer node representations. Then, using the node representations as input, we train a linear classifier to obtain the predictions. The purpose of the linear probing setting is to evaluate the pre-trained model’s ability to generate good node representations. For the fine-tuning setting, we attach a linear classifier to the last layer of the pre-trained encoder and then conduct supervised end-to-end fine-tuning. The purpose of the fine-tuning setting is to evaluate the pre-trained model’s ability to transfer its learned knowledge to unseen graphs and downstream tasks in low-label scenarios. For both settings, we run the experiments 10 times with random seeds.
Results. Table 2 shows the results of our method and baseline methods under the linear probing setting. We have the following observations: (1) PGT outperforms all graph self-supervised learning baselines by a significant margin on both datasets. This demonstrates that PGT can generate more discriminative node representations on unseen graphs. (2) Some pre-trained models exhibit the phenomenon of negative transfer. From the results, especially in Game B, some graph self-supervised learning algorithms perform even worse than the non-pretrained GAT. This suggests that the knowledge learned during pre-training may not effectively transfer to unseen graphs. (3) Comparing the backbone networks, the graph transformer using PPR for sequence sampling significantly outperforms the GAT using ClusterGCN sampling. The reason for this phenomenon could be that ClusterGCN tends to lose over 70-90% of edges when partitioning large graphs, leading to a severe loss of graph structural information.
Table 3 reports the results under the fine-tuning setting. Firstly, it can be observed that all pre-trained models show improvements compared to the non-pretrained model in low-label scenarios. This demonstrates that graph pre-trained models can provide a better initialization weight for downstream tasks, facilitating model fine-tuning under resource-scarce conditions. However, it can be further observed that in scenarios with sufficient labels, some pre-training algorithms fail to consistently bring improvements. One possible reason is that these pre-trained models become overly adapted to specific noise and variations in the training data, leading to a loss of generalization ability. In the vast majority of experimental conditions, PGT consistently outperforms all baseline methods, demonstrating the adaptability of our approach to the widely used pretrain-finetune paradigm in the industry.
5.3. Friend Recall Task
Table 4. Statistics of datasets for Friend Recall Task.

### v4

@@ IV-B Masked Encoding
Unlike most existing masked autoencoder-based graph self-supervised learning frameworks that employ GNNs as encoders [hou2022graphmae, li2022maskgae], in this work, we use the Transformer model as the encoder, which introduces differences in the pipeline.
@@
For a seed node $v_{s}$ in the graph, we sample a sequence of contextual nodes $S_{s}=(v_{s},v_{s,1},v_{s,2},\ldots,v_{s,k})$ of length $k+1$ including seed node $v_{s}$ itself. A subset of this contextual sequence is randomly masked, resulting in a masked set $S^{m}_{s}=\{v_{s,m_{1}},v_{s,m_{2}},\ldots,v_{s,m_{l}}\}$ of length $l$, where $l<k+1$. The remaining unmasked nodes are represented as $S^{u}_{s}=\{v_{s,u_{1}},v_{s,u_{2}},\ldots,v_{s,u_{k-l+1}}\}$. The node feature $\mathbf{x}_{s,u_{j}}\in\mathbb{R}^{1\times d}$ of each unmasked node $v_{s,u_{j}}$ is first mapped into an embedding space using a linear projection from $d$ dimension to $d^{\prime}$ dimension, where $0<j\leq k-l+1$. In addition, we add pre-computed positional encodings to the node embedding:
$\mathbf{h}_{s,u_{j}}=\mathrm{Linear}(\mathbf{x}_{s,u_{j}})+\mathbf{p}_{s,u_{j}}$ (1)
[…]
When using a GNN as an encoder, masked encoding still relies on the complete graph as input. However, when using a transformer as the encoder, we only need to encode a small portion of nodes (e.g., 15-25%) that have not been masked, significantly improving training efficiency.
IV-C Pre-training Tasks
@@
Feature reconstruction. Given a node sequence $S_{s}$ along with its corresponding node features $\mathbf{X}_{s}\in\mathbb{R}^{(k+1)\times d}$, after the masked encoding phase, we obtain representations $\mathbf{H}^{u}_{s}\in\mathbb{R}^{(k-l+1)\times d^{\prime}}$ for the unmasked node sequence. Similar to MIM [he2022masked], the input to the decoder is the full set of tokens consisting of encoded unmasked nodes and masked nodes. We employ a shared learnable token $\mathbf{h}_{[M]}\in\mathbb{R}^{1\times d^{\prime}}$ to initialize the representations of these masked nodes, and this learnable token is continually updated during training. At the same time, we once again add pre-computed positional encodings into the corresponding node embeddings to introduce graph structural information to the decoding phase. The input representation matrix $\mathbf{H}_{s}\in\mathbb{R}^{(k+1)\times d^{\prime}}$ of the decoder can be denoted as:
@@
Then, the decoder reconstructs the input features $\mathbf{X}_{s}$ from the latent embeddings $\mathbf{H}_{s}$. In this task, we use another relatively shallow transformer as a decoder, denoted as $f_{D_{1}}(\cdot)$. The reconstructed features are $\mathbf{Z}_{s}=f_{D_{1}}(\mathbf{H}_{s})$. We adopt scaled cosine error [hou2022graphmae] as the objective function for feature reconstruction:
$\mathcal{L}_{1}=\frac{1}{|S^{m}_{s}|}\sum_{v_{s,i}\in S^{m}_{s}}(1-\frac{\mathbf{x}_{i}^{\top}\mathbf{z}_{i}}{\lVert\mathbf{x}_{i}\lVert\cdot\lVert\mathbf{z}_{i}\lVert})^{\gamma}$ (2)
@@
The feature reconstruction process in our framework operates similarly to filling in missing parts of an image based on visible portions. For each seed node, after identifying its contextual neighborhood and masking approximately 85% of the nodes, the transformer encoder analyzes the few visible nodes to learn their relationships in the graph’s latent space. The decoder then attempts to ”imagine” what the hidden nodes’ features should be, based solely on three sources of information: the processed embeddings from visible nodes, positional information about the masked nodes in the sequence, and a generic learnable token ([M]) indicating a hidden node. By comparing these predictions with the actual features and minimizing the difference, the model learns to understand feature relationships within local neighborhoods, recognize common feature patterns, and develop an implicit understanding of how structural positions correlate with node attributes.
Local graph structure reconstruction. Our intuition behind using PPR to sample contextual node sequences is that selecting the most relevant neighboring nodes can create small and densely connected clusters. PPR can be interpreted within the spectral framework as a localized spectral method [green2021statistical]. Spectral clustering essentially operates in the eigenspace of the graph Laplacian matrix, capturing the low-frequency components that represent tightly-knit communities. [tremblay2020approximating]. Similarly, the embeddings learned by the proposed PGT should ideally reside in a subspace where the local structural properties are maintained. This task aims to capture local graph structure patterns by leveraging the likelihood that nodes within the same sequence exhibit higher similarity, in contrast to nodes from different sequences, which are generally less similar.
@@
$\mathcal{P}$ $=\{(\mathbf{h}_{i,j},\mathbf{h}_{i,k}):i\in\{1,.N\},j,k\in\{1,.M_{i}\},j\neq k\},$
$\mathcal{N}$ $=\{(\mathbf{h}_{i,j},\mathbf{h}_{l,m}):i,l\in\{1,.N\},i\neq l,$
$j\in\{1,.M_{i}\},m\in\{1,.M_{l}\}\}.$
@@
$s_{a,b}=\text{sim}\left(f_{D_{2}}(\mathbf{h}_{a}),f_{D_{2}}(\mathbf{h}_{b})\right)$ (3)
@@
$\mathcal{L}_{2}=-\frac{1}{T}\sum_{(\mathbf{h}_{i,j},\mathbf{h}_{i,k})\in\tilde{\mathcal{P}}}\log\left(\frac{\exp(s_{ij,ik})}{\sum_{(\mathbf{h}_{l,m},\mathbf{h}_{n,o})\in\tilde{\mathcal{P}}\cup\tilde{\mathcal{N}}}\exp(s_{lm,no})}\right).$ (4)
@@
$\mathcal{L}=\mathcal{L}_{1}+\lambda\cdot\mathcal{L}_{2}$ (5)
IV-D Decoder Reuse for Feature Augmentation
Figure 5: Comparison of existing methods and the proposed PGT framework during the inference phase. Our contributions are: a) we use fast approximations of PPR to select informative auxiliary nodes, instead of employing layer-by-layer full neighborhood aggregation. b) We propose a lightweight feature augmentation strategy that requires no additional training. Using the pre-trained encoder and decoder to perform a single forward pass generates reconstructed features, which, when averaged with the original features, serve as the input for downstream tasks.
Under the classic auto-encoder paradigm, only the pre-trained encoder is used in downstream tasks, and the decoder is discarded [he2022masked, hou2022graphmae, hou2023graphmae2, li2022maskgae]. In PGT, we propose a novel approach to utilize the pre-trained decoder for feature augmentation when the pre-trained encoder is used for inference or finetuned for downstream tasks. Given a contextual node sequence $S_{q}$ of a query node $v_{q}$, we generate reconstructed node features by passing the node sequences and their features $\mathbf{X^{\prime}_{q}}\in\mathbb{R}^{(k+1)\times d}$ into the pre-trained encoder and decoder for features reconstruction without any mask:
$\hat{\mathbf{H^{\prime}_{q}}}=f_{E}(\mathrm{Linear}(\mathbf{X^{\prime}_{q}})+\mathbf{P^{\prime}_{q}})$ (6)
$\overline{\mathbf{X^{\prime}_{q}}}=f_{D_{1}}(\hat{\mathbf{H^{\prime}_{q}}}+\mathbf{P^{\prime}_{q}})$ (7)
@@ IV-D Decoder Reuse for Feature Augmentation
Inference. Fig. 5 shows the comparison between existing methods and the proposed PGT framework during the inference phase. The process of utilizing pre-trained decoder for feature augmentation is decoupled from the subsequent fine-tuning of encoder for downstream tasks. After obtaining the newly augmented features, decoders are discarded, and only the encoder is employed for generating node representations or subject to fine-tuning. We take the output of the last layer of encoder as node embeddings:
$\mathbf{H^{\prime}_{q}}=f_{E}(\mathrm{Linear}(\frac{1}{2}(\mathbf{X^{\prime}_{q}}+\overline{\mathbf{X^{\prime}_{q}}}))+\mathbf{P^{\prime}_{q}})$ (8)
@@
V Performance on Public Benchmarks
PGT is designed as a general graph pre-training framework. Before deploying our method on Tencent, we conduct experiments on public benchmarks to validate its effectiveness and broad applicability.
V-A Pre-training
TABLE I: Statistics of public datasets.
Datasets #Nodes #Edges #Features Domain
Cora 2,708 5,429 300 Citation
PubMed 19,717 44,338 300 Citation
ogbn-arxiv 169,343 1,166,243 300 Citation
ogbn-papers100M 111,059,956 1,615,685,872 300 Citation
ogbn-products 2,449,029 61,859,140 300 E-commerce
Wiki-CS 11,701 216,123 300 Web content
FB15K237 14,541 310,116 300 Knowledge
WN18RR 40,943 93,003 300 Knowledge
Dataset. We conduct pre-training on a large paper citation network ogbn-papers100M [hu2021ogb]. In this dataset, each node represents a paper and each directed edge indicates a citation. Each node is associated with its natural language title and abstracts. We leverage word2vec [mikolov2013distributed] to generate unified vector features for all nodes. The statistics are listed in Table I.
Baselines. The baselines are consistent with those on Tencent’s data in Section X. To maximize the performance of the baselines, we adopt the widely used GraphSAINT [zeng2019graphsaint] as the subgraph sampling method in our baselines. For all baseline methods, we employ GAT [velivckovic2017graph] as the backbone network.
V-B In-domain Node Classification Task
TABLE II: Linear probing results on Node Classification Task. We report accuracy(%) for all datasets. No-pretrain represents a random-initialized model without any pre-training.
Methods Cora PubMed ogbn-arxiv
[…]
Improv. vs Best Baseline (%) 3.14% 0.38% 1.35%
Setup. In this study, we adopt Cora [mccallum2000automating], PubMed [sen2008collective] and ogbn-arxiv [hu2021ogb], three popular citation network benchmarks for node classification. In these datasets, each node represents a paper and each directed edge indicates a citation. Each node is associated with its natural language title and abstracts. In alignment with the pre-training graph, we leverage word2vec [mikolov2013distributed] to generate unified vector features for all datasets. The statistics are listed in Table I. For Cora and PubMed, we follow the public data splits as in [velivckovic2018deep, hou2022graphmae, he2023explanations]. For ogbn-arxiv, we follow the official splits in [hu2021ogb]. We select all the self-supervised pre-trained models mentioned earlier as baseline methods. To maximize the performance of the baselines, we employ full neighborhood aggregation in inference. We evaluate our method under linear probing settings.
Results. Table II shows the results of our method and baselines under linear probing setting. We can observe that our proposed PGT framework generally outperforms all baselines across all three datasets. This substantiates that PGT is a general and effective graph pre-training framework, capable of applicability across a diverse range of graph domains. We also observe that some baseline methods manifest severe cases of negative transfer. We hypothesize that a potential reason could be that the citation networks themselves do not possess numerical features in the same feature space. The process of encoding textual features using pre-trained language models may introduce unavoidable noise. The noise in node features adversely impacts the generalization capability of the pre-trained models. Our proposed method for feature augmentation, which utilizes reconstructed features based on masked autoencoders, can be regarded as a form of feature denoising [vincent2008extracting]. During inference, the model applies its learned error-correction ability to the unmasked input, benefiting from the robustness learned during the training phase [vincent2010stacked].
V-C Cross-domain Evaluation on Diverse Tasks
TABLE III: Linear probing results on datasets from diverse domains. We report accuracy (%) for both node classification tasks (ogbn-products, Wiki-CS) and edge classification tasks (FB15K237, WN18RR). No-pretrain represents a random-initialized model without any pre-training.
Methods Node Classification Edge Classification
ogbn-products Wiki-CS FB15K237 WN18RR
[…]
Improv. (%) 4.13% 5.05% 5.65% 7.59%
TABLE IV: Dynamic link prediction results on the time-split ogbn-arxiv dataset. We report MRR (%) and Hits@k (%). Static baselines are trained on the aggregated historical graph.
Method MRR Hits@10 Hits@50
Static Baselines
GCN [kipf2016semi] 32.14 46.81 66.53
GraphSAGE [hamilton2017inductive] 33.59 48.72 68.21
GAT [velivckovic2017graph] 34.72 50.15 70.09
Dynamic Baselines
EvolveGCN [pareja2020evolvegcn] 35.81 51.34 71.82
DGCN [gao2022novel] 36.54 52.01 72.45
VGRNN [hajiramezanali2019variational] 37.92 54.66 74.13
TGN [rossi2020temporal] 41.88 60.29 80.54
PGT-Dynamic (Ours) 45.12 64.77 84.36
Improv. vs Best Baseline (%) 7.74% 7.43% 4.74%
TABLE V: Statistics of datasets for pre-training and Minor Detection Task.
@@ V-C Cross-domain Evaluation on Diverse Tasks
TABLE VI: Linear probing results on Minors Detection Task. We report ROC-AUC for all datasets. No-pretrain represents a random-initialized model without any pre-training.
@@
Improv. vs Best Baseline (%) 4.50% 1.24%
TABLE VII: Results of fine-tuning the pre-trained model with 1%, 5% and 100% labeled training data on Minors Detection Task. We report ROC-AUC for all datasets. No-pretrain represents a random-initialized model without any pre-training.
@@
Improv. vs Best Baseline (%) -0.28% 0.66% 1.50% 1.10% 1.90% 1.30%
Setup. To evaluate the broad applicability of PGT beyond citation networks and gaming scenarios, we conduct additional experiments on four datasets from diverse domains: ogbn-products, Wiki-CS, FB15K237, and WN18RR. These datasets represent varied application scenarios in e-commerce, web content networks, and knowledge graphs. For ogbn-products, each node represents a product sold on Amazon, and edges indicate that the products are frequently purchased together. Wiki-CS contains Wikipedia articles about computer science topics, with edges representing hyperlinks between articles. FB15K237 and WN18RR are knowledge graph datasets where nodes represent entities and edges represent relationships between them. Following our main experimental protocol, we pre-train on ogbn-papers100M and evaluate performance on these target datasets using linear probing. For node classification tasks (ogbn-products and Wiki-CS), we train a linear classifier on top of the frozen node embeddings. For edge classification tasks (FB15K237 and WN18RR), we follow standard practice and concatenate the embeddings of node pairs to train a linear classifier.
Results. Table III shows the results on these diverse datasets. We observe that PGT consistently outperforms all baseline methods across different domains and tasks, with improvements ranging from 4.13% to 7.59% over the best baseline. This demonstrates the strong generalization capability of our pre-training framework beyond citation networks. Notably, on the knowledge graph datasets (FB15K237 and WN18RR), some baseline methods like DGI show significant performance degradation, likely due to the structural differences between the pre-training and target graphs. In contrast, PGT maintains robust performance, suggesting that our approach effectively mitigates negative transfer issues. The superior performance on e-commerce networks (ogbn-products) and web content networks (Wiki-CS) further confirms that PGT’s pre-training strategy captures transferable knowledge applicable to diverse graph structures and domains. These results substantiate our claim that PGT is a general graph pre-training framework with broad applicability across various real-world scenarios beyond our primary application in the gaming industry.
V-D Dynamic Link Prediction Task
To demonstrate the versatility of our framework and its applicability to temporal tasks, we adapt PGT for dynamic link prediction. We introduce PGT-Dynamic, a straightforward yet effective extension of our model. For each graph snapshot at time $t$, we first use the powerful, pre-trained PGT encoder (frozen) to generate high-quality node embeddings $H_{t}$ that capture the intra-snapshot structure. These embeddings are then fed into a Gated Recurrent Unit (GRU) [chung2014empirical], which maintains a hidden state to aggregate temporal information across snapshots. The final, temporally-aware node representation at time $t$ is the output of the GRU. This hybrid approach leverages the rich features from our static pre-training while a lightweight recurrent model captures the temporal dynamics.
Setup. We adapt the ogbn-arxiv dataset for a dynamic link prediction task. Using the publication year of each paper, we create yearly graph snapshots. The task is to predict new citations (links) in a future year based on the graph’s evolution. We use snapshots from before 2017 for training, the 2018 snapshot for validation, and the 2019 snapshot for testing. We compare PGT-Dynamic against both static and dynamic GNN baselines. The static baselines (GCN [kipf2016semi], GraphSAGE [hamilton2017inductive], GAT [velivckovic2017graph]) are trained on an aggregated graph of all training snapshots. The dynamic baselines include EvolveGCN [pareja2020evolvegcn], DGCN [gao2022novel], TGN [rossi2020temporal], and VGRNN [hajiramezanali2019variational]. These are trained sequentially on the snapshots. We evaluate all methods using standard link prediction metrics: Mean Reciprocal Rank (MRR) and Hits@k.
Results. As shown in Table IV, our proposed PGT-Dynamic significantly outperforms all baselines. We observe that the standard static GNNs perform poorly, as they are incapable of capturing the temporal evolution of the graph, confirming the necessity of temporal modeling. Among the specialized dynamic methods, TGN is the strongest baseline. Even so, PGT-Dynamic surpasses TGN by a considerable margin (e.g., a 7.74% relative improvement in MRR).
This result is highly compelling because the dynamic baselines are specifically designed to model temporal evolution from scratch, whereas our adaptation is a simple combination of a static encoder and a GRU. The superior performance of PGT-Dynamic stems from the powerful node representations generated by the pre-trained PGT encoder. These features, learned from the massive ogbn-papers100M dataset, provide a much richer semantic and structural starting point than the randomly initialized features used by the other models. The GRU can then focus on learning the temporal evolution from a high-quality representation space. This experiment strongly suggests that a powerful pre-trained static model can serve as a robust foundation for building effective dynamic models, highlighting the generalization and flexibility of the PGT framework.
VI Deployment in Tencent
VI-A Pre-training
TABLE VIII: Statistics of datasets for Friend Recall Task.